OpenStudy (abbycross167):

Can someone please help me? I don't know how to do it, and I would really like it if someone could show me the steps on how to answer it. . Find the slope-intercept form of the equation that goes through these two points. (1,20) and (8,4.5)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you have to find the slope first

OpenStudy (misty1212):

\[m=\frac{y_2-y_1}{x_2-x_1}\] put \[x_1=1,x+2=8,y_1=20,y_2=4.5\] and compute that fraction

OpenStudy (misty1212):

ooops i means \(x_2=8\)

OpenStudy (abbycross167):

so x2=8 not x+2=8?

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

you should have \[\frac{4.5-20}{8-1}\]

OpenStudy (abbycross167):

ok awesome, sorry if I don't reply fast but I'm writing down the steps

OpenStudy (abbycross167):

may I ask what comes next?

OpenStudy (igreen):

Just simplify what she gave you.

OpenStudy (igreen):

\(\sf \dfrac{4.5-20}{8-1}\)

OpenStudy (abbycross167):

-15.5 7 or 15.5 7?

OpenStudy (benlindquist):

keep the 4.5 - 20 = on the top

OpenStudy (benlindquist):

8 -1 = on the bottom

OpenStudy (igreen):

The first one is correct.

OpenStudy (igreen):

4.5 - 20 will give you a negative number, not a positive.

OpenStudy (benlindquist):

cause u subtract smaller to larger

OpenStudy (abbycross167):

ok then do I divide -15.5 by 7?

OpenStudy (benlindquist):

yes

OpenStudy (abbycross167):

-15.5/7= -2.21428571429?

OpenStudy (benlindquist):

round

OpenStudy (igreen):

Yes, that's correct.

OpenStudy (igreen):

Depending on your answers, you probably have to round.

OpenStudy (abbycross167):

-2.21?

OpenStudy (benlindquist):

-2.2

OpenStudy (abbycross167):

oh ok

OpenStudy (abbycross167):

is that the answer to the question?

OpenStudy (benlindquist):

yes

OpenStudy (abbycross167):

ok great! thank y'all soooo much!! it really helps!!

OpenStudy (benlindquist):

medal please?